HUBBLE SPACE TELESCOPE DAILY REPORT #2230 PERIOD COVERED: 0000Z (UTC) 10/15/98 - 0000Z (UTC) 10/16/98 Daily Status Report as of 289/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Three Sets of WF/PC-2 7713 (WF/P-2 Cycle 7 Supplemental Darks pt3/3) The WF/PC-2 was used to make supplemental dark calibration images. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were completed as planned, and no anomalies were noted. 1.2 Completed Two Sets of STIS/CCD 7948 (CCD Dark and Bias Monitor -- Cont. to Dec. 98) The Space Telescope Imaging Spectrograph (CCD) was used to measure CCD dark current and bias rates for the two supported amplifier settings {GAIN=1 and GAIN=4} in order to gather data for making superdarks and superbiases. The data will be used to track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed Two Sets of NIC/2 7226 (Search for Massive Jupiters) The Near Infrared Camera (Camera Two) was used to make a corographic search for massive planets around nearby, young main sequence stars. Because of the extreme youth of these objects, any low-mass brown dwarf and planetary companions will still be in a higher luminosity phase and thus easily detectable. A dark calibration image was also taken. As described in 2.1 and HSTARs 6718 and 6719, during the second iteration of this proposal, the guide star default to fine lock backup on one FGS only, and the loss of lock and subsequent failures to re-acquire occurred, leaving the take data flag down for most of the iteration. Otherwise, the observations were executed as scheduled, and no further anomalies were noted. 1.4 Completed WF/PC-2 7387 (T-Tauri Star Snapshot Survey) The WF/PC-2 was used to make observations of the variable stars SZ-Cha and SZ-6as part of a T-Tauri class star survey. The observations were executed as scheduled, and no anomalies were noted. 1.5 Completed Two Sets of NIC/1/2/3 7963 (DARKs: Warming Up) The Near Infrared Camera (Cameras One, Two and Three) was used to monitor the detector dark current during the interval between the end of science observations and the exhaustion of cryogen and subsequent warming of the dewar to > 100K. The observations were completed as planned, and no anomalies were noted. 1.6 Completed WF/PC-2 6703 (Weak Radio Galaxies at Z > 2: Imaging a Unique Sample) The WF/PC-2 was used to make observations of the weak radio galaxy GAL-1758+6 which is located at a distance of Z > 2.0. Only the HST can deliver the resolution necessary to investigate the structure of galaxies at redshifts z > 2. As described in yesterday's report and HSTAR 6716, the guide star acquisition for this proposal defaulted to fine lock backup on one FGS only. Otherwise, the observations were completed as scheduled, and no further anomalies were reported. 1.7 Completed Three Sets of NIC/2/3 8082 (NICMOS pure parallel Fall 98) The Near Infrared Camera (Cameras Two and Three) was used to make pure parallel observations while another instrument is prime. As described in yesterday's report and HSTAR 6716, the guide star acquisition for the first iteration of this proposal defaulted to fine lock backup on one FGS only. Otherwise, the observations were executed as scheduled, and no further anomalies were noted. 1.8 Completed Two Sets of STIS/CCD 8062 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation) The Space Telescope Imaging Spectrograph (CCD) was used to make high galactic latitude observations take in parallel with the prime science observations. The observations were completed as planned, and no problems were reported. 1.9 Completed STIS/CCD 8064 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal - Backup) The Space Telescope Imaging Spectrograph (CCD) was used to conduct a non-scripted parallel POMS test. The observations were completed as scheduled, and no problems were reported. 1.10 Completed NIC/2 7894 (Completing A Near-Infrared Search for Very Low Mass Companions to Stars within 10 pc) The Near Infrared Camera (Camera Two) was used to make observations searching for very low mass (VLM) companions to the star BD-08D4352. This search will be the largest, most sensitive, volume-limited search for VLM companions ever undertaken. The observations were completed as planned, and no anomalies were reported. 1.11 Completed NIC/2 7835 (A Search for Superplanets Embedded in Beta Pic & Vega-like Circumstellar Disks) The Near Infrared Camera (Camera Two) was used to make observations of the circumstellar disk around star GL229B in order to search for superplanets. The discovery of a circumstellar disk orbiting Beta Pictoris as well as the discovery of other similar IR-excess Vega-like systems have provided compelling evidence that our solar system, or at least many of the processes which are thought to have created it, may be far from unique. The observations were executed as scheduled, and no anomalies were noted. 1.12 Completed STIS/CCD 7635 (STIS CCD Hot Pixel Annealing) The Space Telescope Imaging Spectrograph (CCD) was used to test the hot pixel annealing procedure. The effectiveness of the CCD hot pixel annealing process is assessed by measuring the dark current behavior before and after annealing and by searching for any window contamination effects. The observations were completed as planned, and no problems were reported. 1.13 Completed NIC/1/2/3 7505 (The Hubble Diagram for Distant Supernovae Measuring Cosmic Deceleration and Global Curvature) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of Super Nova SN1998ESTA in order to produce a precise Hubble diagram from measuring cosmic deceleration and global curvature. HST can make an essential contribution by providing photometry with substantially higher precision and better temporal coverage than can be obtained from the ground. The observations were completed as planned, and no anomalies were reported. 1.14 Completed WF/PC-2 7909 (POMS Test Proposal: WF II Parallel Archive Proposal) The WF/PC-2 was used to make parallel observations as part of the POMS test proposal. This test proposal is designed to simulate scientific plans. The observations were completed as planned, and no problems were reported. 1.15 Completed NIC/1/2/3 7172 ([Fe-II] AND [Si-VI] Line Imaging of Seyfert Galaxies) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of Seyfert Galaxy CIRCINUS to study Fe-II and Si-VI line imaging. The two forbidden lines [Fe-II] and [Si-VI] have very different critical densities and would be expected to originate from very different regions in a Seyfert galaxy nucleus. The observations were completed as planned, and no problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 12 Successful : 12 As documented in HSTAR 6718, the acquisition at 288/123513Z defaulted to fine lock backup on FGS-1 only when FGS-2 exceeded the scan step limit. The proposal described in 1.3 may have been affected. Then, at 288/135452, the same proposal suffered a loss of lock. Subsequent attempts to re-acquire were unsuccessful when the scan step limit was repeatedly exceeded. Scheduled Reacquisitions : 6 Successful : 6 2.2 FHST Updates: Scheduled : 23 Successful : 23 2.3 Operations Notes: As directed by ROP SR-1, the SSR EDAC error counter was cleared twice. Using ROP NS-3, the NSSC-1 status buffer was dumped and reset at 288/1634Z. Per HSTARs 6717, 6720, and 6721, there were three occurrences of NICMOS EDAC single bit errors at 288/1210Z, at 289/0156Z, and at 289/0713Z. In each case, the NICMOS internal log was dumped, and the ground system limit was raised by one. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH